OpenStudy (anonymous):

What expression defines the artmetric series 15+25+35+... for seven terms?

OpenStudy (anonymous):

as we can see 25-15=35-25=10(common difference)its an arithmetic series a,the first term=15,so 7th term,a7=15+(7-1)*10 =15+60 =75

OpenStudy (anonymous):

sum to 7th term, S7 =(7/2)*(a7+a)=(0.5)*(75+15)*7=(0.5)*(630)=315

OpenStudy (anonymous):

So the would be

OpenStudy (anonymous):

Nice job!

OpenStudy (anonymous):

thanks